# Some characterizations of weakly uniformly perfect sets

### v5

@@ 1. Introduction
$I$ $=$ $\frac{1}{2\pi\beta\log\log\frac{1}{\delta(z_{0})}+2\pi\beta\log\left(1+(\log\frac{1}{\delta(z_{0})})^{-1}(\log\frac{\sqrt{C}}{2}-\frac{\beta}{2}\log\log\frac{1}{\delta(z_{0})})\right)-2\pi\log C}$
$+\frac{32C}{3\pi}\left(\log\frac{1}{\delta(z_{0})}-\frac{\beta}{2}\log\log\frac{1}{\delta(z_{0})}+\log\frac{\sqrt{C}}{2}\right)^{-\beta}$
$=$ $\frac{1}{2\pi\beta\log\log\frac{1}{\delta(z_{0})}+O\left(\frac{\log\log\frac{1}{\delta(z_{0})}}{\log\frac{1}{\delta(z_{0})}}\right)}+O\left(\left(\log\frac{1}{\delta(z_{0})}\right)^{-\beta}\right).$
On the other hand, we have $K_{A}(z_{0})\geq K_{\Omega}(z_{0})\geq M\cdot\frac{1}{\delta(z_{0})^{2}\log\log\frac{1}{\delta(z_{0})}}$ for some $M>0$. Thus for each $\beta,C$ and $r_{0}>0$, we find a point $z_{0}\in\Omega$ with $\delta(z_{0})<r_{0}$, such that
$0<M\leq I\cdot\log\log\frac{1}{\delta(z_{0})}.$
Let $C$ be fixed and $r_{0}\to 0$, we obtain
$00<M\leq\frac{1}{2\pi\beta},\,\,\forall\beta>$ 0 .
But it is impossible for $\beta\gg 1$, which implies $(3)$.
@@
We denote the Poincaré metric of a hyperbolic domain $\Omega\subset\mathbb{C}$ by $\rho_{\Omega}(z)|\mathrm{d}z|$, the curvature of which is $-1$. It is obtained by Beardon-Pommerenke [4] that
(3.1) $\rho_{\Omega}(z)\asymp\frac{1}{\delta(z)(\beta_{\Omega}(z)+C)},$
@@
$\beta_{\Omega}(z)=\inf\left\{\left|\log\left|\frac{z-a}{b-a}\right|\right|;a\in\partial\Omega,b\in\partial\Omega,|z-a|=\delta(z)\right\}.$
Let $\partial\Omega$ be $h$-uniformly perfect. For each $w\in\Omega$, take $a\in\partial\Omega$ with $|w-a|=\delta(w)$. If $w$ is sufficiently close to $\partial\Omega$, then there exists a point
$b\in\{z\in\mathbb{C};h(\delta(w))\leq|z-a|\leq\delta(w)\}\cap\partial\Omega.$
@@
$\beta_{\Omega}(w)\leq\left|\log\frac{|w-a|}{|b-a|}\right|\leq\log\frac{\delta(w)}{h(\delta(w))}.$
@@
$\rho_{\Omega}(z)\gtrsim\frac{1}{\delta(z)\log\frac{\delta(z)}{h(\delta(z))}}.$
In particular, if $\partial\Omega$ satisfies $(U)_{1,\alpha}$ or $(U)_{2,\beta}$ , then we have
$\rho_{\Omega}(z)\gtrsim\frac{1}{\delta(z)\log\frac{1}{\delta(z)}}$
@@
$\rho_{\Omega}(z)\gtrsim\frac{1}{\delta(z)\log\log\frac{1}{\delta(z)}},$
@@
The proof of Theorem 1.2 is almost the same as the proof of Theorem 1.1, except that the Bergman kernel is replaced by Poincaré metric. For the completeness, we provide the details here. Recall that for $R=\left\{z\in\mathbb{C};\frac{1}{R}<|z|<R\right\}$, we have
@@
$(1)$ Let $t\in(0,\frac{1}{2}]$. Suppose on the contrary that $\partial\Omega$ does not satisfies condition $(U)_{1,\alpha}$ for each $\alpha>1$. Then for every $C,r_{0}>0,$ there exist $a\in\partial\Omega,r\in(0,r_{0}),$ such that
@@
$T:A\to R:=\left\{z\in\mathbb{C};C^{\frac{1}{2}}r^{\frac{\alpha-1}{2}}<|z|<C^{-\frac{1}{2}}r^{\frac{1-\alpha}{2}}\right\},z\to C^{\frac{1}{2}}r^{-\frac{\alpha+1}{2}}(z-a)$
and $z_{1}=T(z_{0})$, then $|z_{1}|=C^{t-\frac{1}{2}}r^{t(\alpha-1)+\frac{1-\alpha}{2}}$. Hence (3.2) implies
@@
$\leq$ $\frac{\pi}{2\log(C^{-\frac{1}{2}}r^{\frac{1-\alpha}{2}})\delta(z_{0})}\cdot\frac{1}{\cos\left(\frac{\pi}{2}(1-2t)\right)}.$
@@
$\rho_{A}(z_{0})$$\leq$ $\frac{(t(\alpha-1)+1)\pi}{(\alpha-1)\delta(z_{0})}\cdot\frac{1}{\log\frac{1}{\delta(z_{0})}+\log\frac{C^{t}}{2}-\frac{t(\alpha-1)+1}{(\alpha-1)}\log C}\cdot\frac{1}{\cos\left(\frac{\pi}{2}(1-2t)\right)}.$
Notice that $\rho_{A}(z_{0})\geq\rho_{\Omega}(z_{0})\geq\frac{M}{\delta(z_{0})\log\frac{1}{\delta(z_{0})}}$, since $A\subset\Omega$. Thus for each $\alpha>1,t\in(0,\frac{1}{2}]$ and $C,r_{0}>0$, there exist $r\in(0,r_{0})$ and $z_{0}\in\Omega$ with $\delta(z_{0})<r_{0}$, such that
$1<M$$\leq$ $\frac{(t(\alpha-1)+1)\pi}{(\alpha-1)}\cdot\frac{\log\frac{1}{\delta(z_{0})}}{\log\frac{1}{\delta(z_{0})}+\mathrm{const}(t,C)}\cdot\frac{1}{\cos\left(\frac{\pi}{2}\cdot(1-2t)\right)}.$
Let $r_{0}\to 0$, then
@@
Let $\alpha\to\infty$, then
@@
$(2)$ Suppose that for each $\beta>0$, $\partial\Omega$ does not satisfy $(U)_{2,\beta}$. Then for every $C,r_{0}>0,$ there are $a\in\partial\Omega,r\in(0,r_{0})$, such that
@@
Let $A:=\left\{z\in\mathbb{C};Cr\left(\log\frac{1}{r}\right)^{-\beta}<|z-a|<r\right\}$. For $r_{0}<\mathrm{diam}(\Omega)/2$, $\overline{A}\subset\Omega$. Take $z_{0}\in A$ with $|z_{0}-a|=\sqrt{C}r\left(\log\frac{1}{r}\right)^{-\beta/2}$. Then
@@
Then (3.3) implies
$\rho_{A}(z_{0})=\frac{\pi}{2|z_{0}-a|m}\leq\frac{\pi}{2\delta(z_{0})}\cdot\frac{1}{\frac{\beta}{2}\log\log\frac{1}{r}+\log\frac{1}{\sqrt{C}}}.$
@@
$\rho_{A}(z_{0})$$\leq$ $\frac{\pi}{2\delta(z_{0})}\cdot\frac{1}{\frac{\beta}{2}\log\log\left(\frac{1}{\delta(z_{0})}\cdot\frac{\sqrt{C}}{2}\cdot\left(\log\frac{1}{\delta(z_{0})}\right)^{-\beta}\right)+\log\frac{1}{\sqrt{C}}}$
$\leq$ $\frac{\pi}{2\delta(z_{0})}\cdot\frac{1}{\frac{\beta}{2}\log\log\frac{1}{\delta(z_{0})}+O\left(\frac{\frac{\beta}{2}\log\log\frac{1}{\delta(z_{0})}}{\log\frac{1}{\delta(z_{0})}}\right)}.$
Notice that $\rho_{A}(z_{0})\geq\rho_{\Omega}(z_{0})\geq\frac{M}{\delta(z_{0})\log\log\frac{1}{\delta(z_{0})}}$. Thus for every $\beta>0$ and $C,r_{0}>0,$ there exists $z_{0}\in\Omega$ with $\delta(z_{0})<r_{0}$, such that
$0<M\leq\frac{\pi}{2}\cdot\frac{\log\log\frac{1}{\delta(z_{0})}}{\frac{\beta}{2}\log\log\frac{1}{\delta(z_{0})}+O\left(\frac{\frac{\beta}{2}\log\log\frac{1}{\delta(z_{0})}}{\log\frac{1}{\delta(z_{0})}}\right)}.$
Let $r_{0}\to 0$, we have
$00<M\leq\frac{\pi}{\beta},\,\,\,\forall\beta>$ 0 .
But it is impossible if $\beta\gg 1$, which implies $(2)$. ∎
@@
In [30], Sugawa introduced the concept of uniform $LHMD$ condition, which provides a quantitative characterization of the boundary regularity of the Dirichlet problem. It is also equivalent to the $\Delta$-regularity. Let $\Omega$ be a domain in $\mathbb{C}$. For $a\in\partial\Omega$ and $r>0$, $w_{a,r}(z)$ denotes the harmonic measure of $\partial D(a,r)\cap\Omega$ relative to $D(a,r)\cap\Omega$. In this paper, $\Omega\subset\mathbb{C}$ is said to satisfy the property $(LHMD)_{1,\gamma}$ or $(LHMD)_{2,\eta}$, if there exist some constant $C,r_{0}>0$ such that
@@
for every $a\in\partial\Omega,r\in(0,r_{0})$ and $z\in D(x,r)\cap\Omega$, respectively.
@@
$(1)$ If $\Omega$ satisfies $(LHMD)_{1,\gamma}$ for some $\gamma>0$, then there exists some $\alpha>1$ such that $\Omega$ satisfies $(\Delta)_{1,\alpha}$.
$(2)$ If $\Omega$ satisfies $(LHMD)_{2,\eta}$ for some $\eta>0$, then there exists some $\beta>0$ such that $\Omega$ satisfies $(\Delta)_{2,\beta}$.
@@
$(1)$The condition $(LHMD)_{1,\gamma}$ implies that there exist some constant $C,r_{0}>0$ such that for every $a\in\partial\Omega$ and $r\in(0,r_{0})$ we have
@@
Let $h(r)=r^{\alpha}$ for some $\alpha>1$, then we have
@@
For $\beta\gg 1$, $Ce^{-\eta\beta}<1$. Thus there exists a constant $r_{1}\in(0,r_{0})$ such that $Ce^{\eta I(r)}<1$ for all $0<r<r_{0}$. Let $\varepsilon=1-I(r)$. Then $\Omega$ satisfies the condition $(\Delta)_{2,\beta}$.
@@
$(1)$ If $\Omega$ satisfies $(\Delta)_{1,\alpha}$ for some $\alpha>1$, then there exists some $\alpha^{\prime}>\alpha$ such that $\partial\Omega$ satisfies $(U)_{1,\alpha^{\prime}}$.
$(2)$ If $\Omega$ satisfies $(\Delta)_{2,\beta}$ for some $\beta>0$, then there exists some $\beta^{\prime}>\beta$ such that $\partial\Omega$ satisfies $(U)_{2,\beta^{\prime}}$.
@@
$(1)$ The condition $(\Delta)_{1,\alpha}$ means that there exist some positive constant $\varepsilon,C$ and $r_{0}$, such that for every $a\in\partial\Omega$ and $r\in(0,r_{0})$ we have
@@
We show $(1)$ by contradiction. Suppose on the contrary that for every $\alpha^{\prime}>\alpha,C^{\prime}>0$ and $r_{1}>0$, there exist $a\in\partial\Omega$ and $r\in(0,r_{1})$, such that
@@
$\phi(z):=\frac{\log\frac{|z-a|}{C^{\prime}r^{\alpha^{\prime}}}}{\log\frac{r}{C^{\prime}r^{\alpha^{\prime}}}}\$
@@
$(2)$ The condition $(\Delta)_{2,\beta}$ means that there exist some positive constant $\varepsilon,C$ and $r_{0}$, such that for every $a\in\partial\Omega$ and $r\in(0,r_{0})$ we have
@@
Then for $r<r_{1}<\frac{\mathrm{diam}(\Omega)}{2}$ we have
@@
$\phi(z):=\frac{\log\frac{|z-a|}{C^{\prime}r\left(\log\frac{1}{r}\right)^{-\beta^{\prime}}}}{\log\frac{r}{C^{\prime}r\left(\log\frac{1}{r}\right)^{-\beta^{\prime}}}}.\$
@@
which is a contradiction as $\beta^{\prime}\to+\infty$.
To finish the proof of Theorem 1.3, we need the following result from [11].
@@
$(1)$ If $\Omega$ satisfies $(U)_{1,\alpha}$ for some $\alpha\in(1,2)$, then there exists some $\gamma>0$ such that $\Omega$ satisfies $(LHMD)_{1,\gamma}$.
$(2)$ If $\Omega$ satisfies $(U)_{2,\beta}$ for some $\beta>0$, then there exists some $\eta>0$ such that $\Omega$ satisfies $(LHMD)_{2,\eta}$.
@@
$(1)$ Theorem 1.3 of [34] implies that there exist $C,r_{0}>0$ such that
@@
Let $\kappa\in\left(0,\frac{1}{16}\right)$ and $r_{1}\in\left(0,\frac{r_{0}}{\kappa}\right)$ be fixed. For every $a\in\partial\Omega$ and $r\in(0,r_{1})$, $w_{a,r}$ is a harmonic function on $\Omega_{r}:=\Omega\cap D(a,r)$, and $w_{a,r}(z)=0$ n.e. on $\partial\Omega_{r}\cap D\left(0,\frac{r}{2}\right)$. Then Theorem 4.3 implies for $z\in\Omega_{r}$ with $|z-a|<\frac{\kappa r}{2}$
@@
where $\gamma$ and $C_{1}$ depends on $\kappa,r_{1},\alpha,C$. When $\frac{\kappa r}{2}\leq|z-a|\leq r$, $w_{a,r}\leq 1$, and
@@
where $C_{2}$ depends on $\kappa,r_{1},\alpha,C$. Thus, there exists a $C_{3}>0$ such that
@@
$(2)$ Theorem 1.3 of [34] implies that there exist $C,r_{0}>0$ such that
@@
We use Theorem 4.3 again to see that for $z\in\Omega_{r}$ with $|z-a|<\frac{\kappa r}{2}$
@@
where $\eta$ depends on $\kappa,r_{1},\alpha,C$. Notice that
@@
where $C_{2}$ depends on $\kappa,r_{1},\alpha,C$. Thus, there exists a $C_{3}>0$ such that
@@
Let $g:(0,+\infty)\to(0,+\infty)$ be an increasing function, such that $\lim_{t\to 0}g(t)=0$. Let $E$ be a closed set in $\mathbb{C}$. Recall the concepts of Hausdorff content and Hausdorff measure:
$\Lambda_{g}(E):=\inf\left\{\sum_{k}g(\mathrm{diam}(B_{k}));\{B_{k}\}\text{is a countable covering of $E$ composed of closed discs}\right\},$
$\mathcal{H}_{g}^{\varepsilon}(E):=\inf\left\{\sum_{k}g(\mathrm{diam}(B_{k}));\{B_{k}\}\text{is a countable covering of $E$ such that $diam(B_{k})<\varepsilon$}\right\},$
@@
Let $E\subset\mathbb{C}$ be $h$-uniformly perfect. Then there exists $r_{0}>0$, such that for every $a\in E,r\in(0,r_{0})$ and $\tilde{c}\in(0,\frac{1}{2})$, there are closed discs $B_{1}$ and $B_{2}$, such that
@@
Let $B_{1}=\overline{B}\left(a,\tilde{c}h\left(\frac{r}{2}\right)\right)\subset B$. Since $E$ is $h$-uniformly perfect, there exists a constant $r_{0}>0$, such that
@@
Applying Lemma 5.1 inductively, we select a sequence $\{B_{i_{1},\cdots,i_{k}}\}_{k\in\mathbb{N},i_{j}\in I},I=\{1,2\}$, such that
@@
We equip $I$ and $I^{\mathbb{N}}$ the discrete and product topologies. Obviously, $\mathcal{C}\subset\mathbb{C}$ is Hausdorff, and Tychonoff Theorem implies that $I^{\mathbb{N}}$ is compact. It is also easy to verify that $f$ is a continuous bijection, so $f$ is a homeomorphic mapping. We know that $I^{\mathbb{N}}$ has a standard Bernoulli measure $\nu$ such that $\nu([i_{1},\cdots,i_{k}])=2^{-k}$ for each $[i_{1},\cdots,i_{k}]=\{(j_{l})_{l\in\mathbb{N}};j_{1}=i_{1},\cdots,j_{k}=i_{k}\}$. Then $\mu:=f_{*}\nu$ is a probability measure on $\mathcal{C}$, such that

### v6

@@ 1. Introduction
$I$ $=$ $\frac{1}{2\pi\beta\log\log\frac{1}{\delta_{\Omega}(z_{0})}+2\pi\beta\log\left(1+(\log\frac{1}{\delta_{\Omega}(z_{0})})^{-1}(\log\frac{\sqrt{C}}{2}-\frac{\beta}{2}\log\log\frac{1}{\delta_{\Omega}(z_{0})})\right)-2\pi\log C}$
$+\frac{32C}{3\pi}\left(\log\frac{1}{\delta_{\Omega}(z_{0})}-\frac{\beta}{2}\log\log\frac{1}{\delta_{\Omega}(z_{0})}+\log\frac{\sqrt{C}}{2}\right)^{-\beta}$
$=$ $\frac{1}{2\pi\beta\log\log\frac{1}{\delta_{\Omega}(z_{0})}+O\left(\frac{\log\log\frac{1}{\delta_{\Omega}(z_{0})}}{\log\frac{1}{\delta_{\Omega}(z_{0})}}\right)}+O\left(\left(\log\frac{1}{\delta_{\Omega}(z_{0})}\right)^{-\beta}\right).$
On the other hand, we have $K_{A}(z_{0})\geq K_{\Omega}(z_{0})\geq M\cdot\frac{1}{\delta_{\Omega}(z_{0})^{2}\log\log\frac{1}{\delta_{\Omega}(z_{0})}}$ for some $M>0$. Thus, for each $\beta,C$ and $r_{0}>0$, we can find a point $z_{0}\in\Omega$ with $\delta_{\Omega}(z_{0})<r_{0}$, such that
$0<M\leq I\cdot\log\log\frac{1}{\delta_{\Omega}(z_{0})}.$
Let $C$ be fixed and $r_{0}\to 0$. Then we obtain
$0<M\leq\frac{1}{2\pi\beta},\,\,\forall\beta>0.$
But it is impossible for $\beta\gg 1$, which completes the proof of $(3)$.
@@
We denote the Poincaré metric of a hyperbolic domain $\Omega\subset\mathbb{C}$ by $\rho_{\Omega}(z)|\mathrm{d}z|$, whose curvature is 1-1. It is obtained by Beardon-Pommerenke [4] that
(3.1) $\rho_{\Omega}(z)\asymp\frac{1}{\delta_{\Omega}(z)(\beta_{\Omega}(z)+C)},$
@@
$\beta_{\Omega}(z)=\inf\left\{\left|\log\left|\frac{z-a}{b-a}\right|\right|;a\in\partial\Omega,b\in\partial\Omega,|z-a|=\delta_{\Omega}(z)\right\}.$
Let $\partial\Omega$ be $h$-uniformly perfect. For each $w\in\Omega$, choose a point $a\in\partial\Omega$ with $|w-a|=\delta_{\Omega}(w)$. If $w$ is sufficiently close to $\partial\Omega$, then there exists a point
$b\in\{z\in\mathbb{C};h(\delta_{\Omega}(w))\leq|z-a|\leq\delta_{\Omega}(w)\}\cap\partial\Omega.$
@@
$\beta_{\Omega}(w)\leq\left|\log\frac{|w-a|}{|b-a|}\right|\leq\log\frac{\delta_{\Omega}(w)}{h(\delta_{\Omega}(w))}.$
@@
$\rho_{\Omega}(z)\gtrsim\frac{1}{\delta_{\Omega}(z)\log\frac{\delta_{\Omega}(z)}{h(\delta_{\Omega}(z))}}.$
In particular, if $\partial\Omega$ satisfies $(U)_{1,\alpha}$ or $(U)_{2,\beta}$ , then we obtain
$\rho_{\Omega}(z)\gtrsim\frac{1}{\delta_{\Omega}(z)\log\frac{1}{\delta_{\Omega}(z)}}$
@@
$\rho_{\Omega}(z)\gtrsim\frac{1}{\delta_{\Omega}(z)\log\log\frac{1}{\delta_{\Omega}(z)}},$
@@
The proof of Theorem 1.2 is almost identical to that of Theorem 1.1, except that the Bergman kernel is replaced by Poincaré metric. For completeness, we provide the details here. Recall that for $R=\left\{z\in\mathbb{C};\frac{1}{R}<|z|<R\right\}$, we have
@@
$(1)$ Let $t\in(0,\frac{1}{2}]$. Suppose on the contrary that $\partial\Omega$ does not satisfies condition $(U)_{1,\alpha}$ for every $\alpha>1$. Then, for every $C,r_{0}>0,$ there exist a point $a\in\partial\Omega$ and a radius $r\in(0,r_{0})$ such that
@@
$T:A\to R:=\left\{z\in\mathbb{C};C^{\frac{1}{2}}r^{\frac{\alpha-1}{2}}<|z|<C^{-\frac{1}{2}}r^{\frac{1-\alpha}{2}}\right\},z\to C^{-\frac{1}{2}}r^{-\frac{\alpha+1}{2}}(z-a)$
and $z_{1}=T(z_{0})$. Then $|z_{1}|=C^{t-\frac{1}{2}}r^{t(\alpha-1)+\frac{1-\alpha}{2}}$. Hence, (3.2) implies
@@
$\leq$ $\frac{\pi}{2\log(C^{-\frac{1}{2}}r^{\frac{1-\alpha}{2}})\delta_{\Omega}(z_{0})}\cdot\frac{1}{\cos\left(\frac{\pi}{2}(1-2t)\right)}.$
@@
$\rho_{A}(z_{0})$ $\leq$ $\frac{(t(\alpha-1)+1)\pi}{(\alpha-1)\delta_{\Omega}(z_{0})}\cdot\frac{1}{\log\frac{1}{\delta_{\Omega}(z_{0})}+\log\frac{C^{t}}{2}-\frac{t(\alpha-1)+1}{(\alpha-1)}\log C}\cdot\frac{1}{\cos\left(\frac{\pi}{2}(1-2t)\right)}.$
Notice that $\rho_{A}(z_{0})\geq\rho_{\Omega}(z_{0})\geq\frac{M}{\delta_{\Omega}(z_{0})\log\frac{1}{\delta_{\Omega}(z_{0})}}$, since $A\subset\Omega$. Thus, for each $\alpha>1,t\in(0,\frac{1}{2}]$, and $C,r_{0}>0$, there exist $r\in(0,r_{0})$ and $z_{0}\in\Omega$ with $\delta_{\Omega}(z_{0})<r_{0}$, such that
$1<M$ $\leq$ $\frac{(t(\alpha-1)+1)\pi}{(\alpha-1)}\cdot\frac{\log\frac{1}{\delta_{\Omega}(z_{0})}}{\log\frac{1}{\delta_{\Omega}(z_{0})}+\mathrm{const}(t,C,\alpha)}\cdot\frac{1}{\cos\left(\frac{\pi}{2}\cdot(1-2t)\right)}.$
Letting $r_{0}\to 0$, we obtain
@@
Then letting $\alpha\to\infty$, we conclude
@@
$(2)$ Suppose that for every $\beta>0$, $\partial\Omega$ does not satisfy condition $(U)_{2,\beta}$. Then, for every $C,r_{0}>0,$ there exist a point $a\in\partial\Omega$ and a radius $r\in(0,r_{0})$ such that
@@
Let $A:=\left\{z\in\mathbb{C};Cr\left(\log\frac{1}{r}\right)^{-\beta}<|z-a|<r\right\}$. For $r_{0}<\mathrm{diam}(\Omega)/2$, we have $\overline{A}\subset\Omega$. Take $z_{0}\in A$ with $|z_{0}-a|=\sqrt{C}r\left(\log\frac{1}{r}\right)^{-\beta/2}$. Then
@@
Then, (3.3) implies
$\rho_{A}(z_{0})=\frac{\pi}{2|z_{0}-a|m}\leq\frac{\pi}{2\delta_{\Omega}(z_{0})}\cdot\frac{1}{\frac{\beta}{2}\log\log\frac{1}{r}+\log\frac{1}{\sqrt{C}}}.$
@@
$\rho_{A}(z_{0})$ $\leq$ $\frac{\pi}{2\delta_{\Omega}(z_{0})}\cdot\frac{1}{\frac{\beta}{2}\log\log\left(\frac{1}{\delta_{\Omega}(z_{0})}\cdot\frac{\sqrt{C}}{2}\cdot\left(\log\frac{1}{\delta_{\Omega}(z_{0})}\right)^{-\beta}\right)+\log\frac{1}{\sqrt{C}}}$
$\leq$ $\frac{\pi}{2\delta_{\Omega}(z_{0})}\cdot\frac{1}{\frac{\beta}{2}\log\log\frac{1}{\delta_{\Omega}(z_{0})}+O\left(\frac{\frac{\beta}{2}\log\log\frac{1}{\delta_{\Omega}(z_{0})}}{\log\frac{1}{\delta_{\Omega}(z_{0})}}\right)}.$
Notice that $\rho_{A}(z_{0})\geq\rho_{\Omega}(z_{0})\geq\frac{M}{\delta_{\Omega}(z_{0})\log\log\frac{1}{\delta_{\Omega}(z_{0})}}$. Thus, for every $\beta>0$ and $C,r_{0}>0,$ there exists $z_{0}\in\Omega$ with $\delta_{\Omega}(z_{0})<r_{0}$, such that
$0<M\leq\frac{\pi}{2}\cdot\frac{\log\log\frac{1}{\delta_{\Omega}(z_{0})}}{\frac{\beta}{2}\log\log\frac{1}{\delta_{\Omega}(z_{0})}+O\left(\frac{\frac{\beta}{2}\log\log\frac{1}{\delta_{\Omega}(z_{0})}}{\log\frac{1}{\delta_{\Omega}(z_{0})}}\right)}.$
Letting $r_{0}\to 0$, we obtain
$0<M\leq\frac{\pi}{\beta},\,\,\,\forall\beta>0.$
But it is impossible if $\beta\gg 1$, which completes the proof of $(2)$. ∎
@@
In [30], Sugawa introduced the concept of the uniform $LHMD$ condition, which provides a quantitative characterization of the boundary regularity of the Dirichlet problem. This condition is also equivalent to the uniform $\Delta$-regularity. Let $\Omega$ be a domain in $\mathbb{C}$. For $a\in\partial\Omega$ and $r>0$, $w_{a,r}(z)$ denotes the harmonic measure of $\partial D(a,r)\cap\Omega$ relative to $D(a,r)\cap\Omega$. In this paper, we say that a domain $\Omega\subset\mathbb{C}$ satisfies the property $(LHMD)_{1,\gamma}$ or $(LHMD)_{2,\eta}$ if there exist constant $C,r_{0}>0$ such that
@@
for every $a\in\partial\Omega,r\in(0,r_{0})$, and $z\in D(x,r)\cap\Omega$, respectively.
@@
$(1)$ If $\Omega$ satisfies $(LHMD)_{1,\gamma}$ for some $\gamma>0$, then there exists $\alpha>1$ such that $\Omega$ satisfies $(\Delta)_{1,\alpha}$.
$(2)$ If $\Omega$ satisfies $(LHMD)_{2,\eta}$ for some $\eta>0$, then there exists $\beta>0$ such that $\Omega$ satisfies $(\Delta)_{2,\beta}$.
@@
(1)(1)The condition $(LHMD)_{1,\gamma}$ implies that there exist constant $C,r_{0}>0$ such that for every $a\in\partial\Omega$ and $r\in(0,r_{0})$, we have
@@
Let $h(r)=r^{\alpha}$ for some $\alpha>1$. Then we have
@@
Notice that for $\beta\gg 1$, $Ce^{-\frac{\eta\beta}{2}}<1$. Thus, there exists a constant $r_{1}\in(0,r_{0})$ such that
$Ce^{\eta I(r)}<Ce^{-\frac{\eta\beta}{2}}<1$
for all $0<r<r_{1}$. Let $\varepsilon=1-Ce^{-\frac{\eta\beta}{2}}$. Then $\Omega$ satisfies the condition $(\Delta)_{2,\beta}$.
@@
$(1)$ If $\Omega$ satisfies $(\Delta)_{1,\alpha}$ for some $\alpha>1$, then there exists $\alpha^{\prime}>\alpha$ such that $\partial\Omega$ satisfies $(U)_{1,\alpha^{\prime}}$.
$(2)$ If $\Omega$ satisfies $(\Delta)_{2,\beta}$ for some $\beta>0$, then there exists $\beta^{\prime}>\beta$ such that $\partial\Omega$ satisfies $(U)_{2,\beta^{\prime}}$.
@@
$(1)$ The condition $(\Delta)_{1,\alpha}$ means that there exist positive constant $\varepsilon,C$ and $r_{0}$ such that, for every $a\in\partial\Omega$ and $r\in(0,r_{0})$, we have
@@
We show $(1)$ by contradiction. Suppose on the contrary that for every $\alpha^{\prime}>\alpha,C^{\prime}>0$ and $r_{1}>0$, there exist $a\in\partial\Omega$ and $r\in(0,r_{1})$ such that
@@
$\phi(z):=\frac{\log\frac{|z-a|}{C^{\prime}r^{\alpha^{\prime}}}}{\log\frac{r}{C^{\prime}r^{\alpha^{\prime}}}}\\$
@@
$(2)$ The condition $(\Delta)_{2,\beta}$ means that there exist some positive constants $\varepsilon,C$ and $r_{0}$, such that for every $a\in\partial\Omega$ and $r\in(0,r_{0})$, we have
@@
Then for $r<r_{1}<\frac{\mathrm{diam}(\Omega)}{2}$, we have
@@
$\phi(z):=\frac{\log\frac{|z-a|}{C^{\prime}r\left(\log\frac{1}{r}\right)^{-\beta^{\prime}}}}{\log\frac{r}{C^{\prime}r\left(\log\frac{1}{r}\right)^{-\beta^{\prime}}}}.\\$
@@
which leads to a contradiction as $\beta^{\prime}\to+\infty$.
To complete the proof of Theorem 1.3, we need the following result from [11].
@@
$(1)$ If $\Omega$ satisfies $(U)_{1,\alpha}$ for some $\alpha\in(1,2)$, then there exists $\gamma>0$ such that $\Omega$ satisfies $(LHMD)_{1,\gamma}$.
$(2)$ If $\Omega$ satisfies $(U)_{2,\beta}$ for some $\beta>0$, then there exists $\eta>0$ such that $\Omega$ satisfies $(LHMD)_{2,\eta}$.
@@
$(1)$ Theorem 1.3 of [34] implies that there exist some constants $C,r_{0}>0$ such that
@@
Let $\kappa\in\left(0,\frac{1}{16}\right)$ and $r_{1}\in\left(0,\frac{r_{0}}{\kappa}\right)$ be fixed. For every $a\in\partial\Omega$ and $r\in(0,r_{1})$, $w_{a,r}$ is a harmonic function on $\Omega_{r}:=\Omega\cap D(a,r)$, and $w_{a,r}(z)=0$ n.e. on $\partial\Omega_{r}\cap D\left(0,\frac{r}{2}\right)$. Then Theorem 4.3 implies that for $z\in\Omega_{r}$ with $|z-a|<\frac{\kappa r}{2}$, we have
@@
where $\gamma$ and $C_{1}$ depend on $\kappa,r_{1},\alpha,C$. When $\frac{\kappa r}{2}\leq|z-a|\leq r$, we have $w_{a,r}\leq 1$, and
@@
where $C_{2}$ depends on $\kappa,r_{1},\alpha$ and $C$. Thus, there exists a constant $C_{3}>0$ such that
@@
$(2)$ Theorem 1.3 of [34] implies that there exist constants $C,r_{0}>0$ such that
@@
We use Theorem 4.3 again to deduce that for $z\in\Omega_{r}$ with $|z-a|<\frac{\kappa r}{2}$
@@
where $\eta$ depends on $\kappa,r_{1},\alpha$ and $C$. Notice that
@@
where $C_{2}$ depends on $\kappa,r_{1},\alpha,C$. Thus, there exists a constant $C_{3}>0$ such that
@@
Let $g:(0,+\infty)\to(0,+\infty)$ be an increasing function such that $\lim_{t\to 0}g(t)=0$. Let $E$ be a closed set in $\mathbb{C}$. Recall the concepts of Hausdorff content and Hausdorff measure:
[…]
@@
Let $E\subset\mathbb{C}$ be $h$-uniformly perfect. Then there exists $r_{0}>0$ such that for every $a\in E,r\in(0,r_{0})$, and $\tilde{c}\in(0,\frac{1}{2})$, there are closed discs $B_{1}$ and $B_{2}$, such that
@@
Let $B_{1}=\overline{B}\left(a,\tilde{c}h\left(\frac{r}{2}\right)\right)\subset B$. Since $E$ is $h$-uniformly perfect, there exists a constant $r_{0}>0$ such that
@@
Applying Lemma 5.1 inductively, we select a sequence $\{B_{i_{1},\cdots,i_{k}}\}_{k\in\mathbb{N},i_{j}\in I},I=\{1,2\}$ such that:
@@
We equip $I$ and $I^{\mathbb{N}}$ the discrete and product topologies, respectively. Obviously, $\mathcal{C}\subset\mathbb{C}$ is Hausdorff, and Tychonoff Theorem implies that $I^{\mathbb{N}}$ is compact. It is also easy to verify that $f$ is a continuous bijection, so $f$ is a homeomorphism. We know that $I^{\mathbb{N}}$ carries a standard Bernoulli measure $\nu$ such that $\nu([i_{1},\cdots,i_{k}])=2^{-k}$, for each $i_{1},\cdots,i_{k}]=\{(j_{l})_{l\in\mathbb{N}};j_{1}=i_{1},\cdots,j_{k}=i_{k}\}$. Then $\mu:=f_{*}\nu$ is a probability measure on $\mathcal{C}$ such that